# Constraint-Driven Multi-USV Coverage Path Generation for Aquatic Environmental Monitoring

### v1

@@ IV Simulation with Ideal Mathematical Model
The time evolution of the functions $b^{1}_{\mathrm{e},1}$ (red) and $b^{1}_{\mathrm{e},2}$ (blue) are shown in Fig. 7. We observe from the figure that $b^{1}_{\mathrm{e},i}\geq 0$ is satisfied at almost all times. This implies that the present path generator almost certifies the performance guarantee. However, the constraint is violated only at around $160$\mathrm{s}$$. In order to further investigate the violation, we show the time evolution of the functions $b^{k}_{\mathrm{e},1}\ (k=2,3,4,5)$ for USV 1 in Fig. 7. We see from this figure that $b^{5}_{\mathrm{e},1}$ is close to the lower limit and the constraint associated with $b^{5}_{\mathrm{e},1}$ gets active during the period. We conclude from these figures that the constraint violations in Fig. 7 are caused by the relaxation of the performance constraints to soft constraints in (46).
From these results, we conclude that the present path generator works as expected, at least for an ideal mathematical model (1).
@@ V-C Collision Avoidance against Walls
Figure 15: Snapshots of the experiment at each time. The left side shows moving two real USVs on the experiment pool. The red and green circular paths on the right side, indicate paths corresponding to USVs 1 and 2. White curves show the trajectories of the two USVs after $t=0$.
@@
$\mu(p)=((p-c)\odot(p-c))^{\top}P((p-c)\odot(p-c)),$
where $\odot$ represents the Hadamard product, $c$ is the center of the pool, and $P\in\mathbb{R}^{2\times 2}$ is a positive definite matrix that scales and rotates the 4-norm ball. Define candidates of the control barrier functions as
@@
$\mu_{\rm d}(p)=$ $-4(\mathbf{1}_{2}\odot(p-c))^{\top}P((p-c)\odot(p-c))$
@@ V-D Experimental Verification
The experimental results are further verified through three types of time responses in Figs. 1618. First, Fig. 16 shows the evolution of the functions $b^{1}_{\mathrm{c},1}$ (red) and $b^{1}_{\mathrm{c},2}$ (blue) on the quality of the paths. We see from this figure that satisfaction of the constraints is more challenging than the simulation of Fig. 7. The USVs occasionally fail to follow the paths determined by solving (33) due to the joint effect of the added wall avoidance behavior and uncertainties in the real physical world. However, we observe that, as expected, the present controller pushes up the functions $b^{1}_{\mathrm{c},i}$ whenever they become negative, avoiding violations for long periods of time.
@@
As for the wall collision avoidance, Fig. 17 shows the time evolution of $b_{i}^{\rm right}$ and $b_{i}^{\rm left}$ for USVs $i=1,2$. We see from these figures that the functions $b_{i}^{\rm right}$ (red) and $b_{i}^{\rm left}$ (blue) $\ (i=1,2)$ are kept positive for almost all time. Only $b_{2}^{\rm left}$ violates the constraint at around $180$\mathrm{s}$$ due to the relaxation of the associated constraint (58c) and various uncertainties of the environment and the system model. Nevertheless, we see from the movie that the collisions with the walls actually do not happen throughout the experiment owing to the margin set in the definition of the environment. The movie also shows that the USVs successfully take avoidance actions when they are close to the wall, while giving up the path following.
@@ VI Conclusion
Future work should be directed towards an extension to a more generic class of paths and field experiments.
@@ Appendix A Quadratic Program for Circular Path Generation
Given a turning direction $\mathrm{X}_{i}\in\mathcal{X}_{i}$, we have
@@
$\mathcal{X}_{\mathrm{c},i}^{\epsilon}(r_{i})=\left\{\mathrm{X}_{i}\in\mathcal{X}_{i}\left|\ \left|\left(\max_{\mathrm{X}_{i}\in\mathcal{X}_{i}}I^{{\rm X}_{i}}_{\mathrm{c},i}(r_{i})\right)-I^{{\rm X}_{i}}_{\mathrm{c},i}(r_{i})\right|\leq\epsilon\right.\right\},$ (60)
@@ Appendix B Quadratic Program for Elliptic Path Generation
$\mathcal{X}_{\mathrm{e},i}^{\epsilon}(s_{i})$ =$=\left\{\mathrm{X}_{i}\in\mathcal{X}_{i}\left|\ \left|\left(\max_{\mathrm{X}_{i}\in\mathcal{X}_{i}}I^{{\rm X}_{i}}_{\mathrm{e},i}(s_{i})\right)-I^{{\rm X}_{i}}_{\mathrm{e},i}(s_{i})\right|\leq\epsilon\right.\right\},$

### v2

@@ IV Simulation with Ideal Mathematical Model
The time evolution of the functions $b^{1}_{\mathrm{e},1}$ (red) and $b^{1}_{\mathrm{e},2}$ (blue) are shown in Fig. 7. We observe from the figure that $b^{1}_{\mathrm{e},i}\geq 0$ is satisfied at almost all times. This implies that the present path generator almost certifies the performance guarantee. However, the constraint is violated only at around $160$\mathrm{s}$$. In order to further investigate the violation, we show the time evolution of the functions $b^{k}_{\mathrm{e},1}\ (k=2,3,4,5)$ for USV 1 in Fig. 7. We see from this figure that $b^{5}_{\mathrm{e},1}$ is close to the lower limit and the constraint associated with $b^{5}_{\mathrm{e},1}$ gets active during the period. We conclude from these figures that the constraint violations in Fig. 7 are caused by the relaxation of the performance constraints to soft constraints in (46). Note that the constraint on $b^{5}_{\mathrm{e},1}$ is not only violated by the initial state but also slightly violated at around 160s. The latter violation was caused by the unavoidable discretization in the path computation. It is however well-known that the constraint-based control inherently drives the states to satisfy the constraints even if the state gets out of the safe set for some reason. This is why these temporal constraint violations do not cause any serious problem.
From these results, we conclude that the present path generator works as expected, at least for an ideal mathematical model (1). Finally, in order to show that the proposed path generator works even for a larger fleet of USVs, a simulation video with 6 USVs is uploaded to https://youtu.be/oS3x3O_IQB8 for reference.
@@ V-C Collision Avoidance against Walls
Figure 15: Snapshots of the experiment at each time. The left side shows moving two real USVs on the experiment pool. The red and green circular paths on the right side, indicate paths corresponding to USVs 1 and 2. White curves show the trajectories of the two USVs after $t=0$\mathrm{s}$$.
@@
$\mu(p)=((p-o)\odot(p-o))^{\top}P((p-o)\odot(p-o)),$ (54)
where $\odot$ represents the Hadamard product, $o$ is the center of the pool, and $P\in\mathbb{R}^{2\times 2}$ is a positive definite matrix that scales and rotates the 4-norm ball. Define candidates of the control barrier functions as
@@
$\mu_{\rm d}(p)=$ $-4(\mathbf{1}_{2}\odot(p-o))^{\top}P((p-o)\odot(p-o))$
@@ V-D Experimental Verification
The experimental results are further verified through three types of time responses in Figs. 1618. First, Fig. 16 shows the evolution of the functions $b^{1}_{\mathrm{c},1}$ (red) and $b^{1}_{\mathrm{c},2}$ (blue) on the quality of the paths. We see from this figure that satisfaction of the constraints is more challenging than the simulation of Fig. 7. The USVs occasionally fail to follow the paths determined by solving (33) due to the joint effect of the added wall avoidance behavior and uncertainties in the real physical world. However, we observe that, as expected, the present controller pushes up the functions $b^{1}_{\mathrm{c},i}$ whenever they become negative, avoiding violations for long periods of time. Also, the path for USV $1$ seems inefficient during the period from $230$\mathrm{s}$$ to $250$\mathrm{s}$$ in the experiment, but the path certainly meets the performance constraint.
@@
As for the wall collision avoidance, Fig. 17 shows the time evolution of $b_{i}^{\rm right}$ and $b_{i}^{\rm left}$ for USVs $i=1,2$. We see from these figures that the functions $b_{i}^{\rm right}$ (red) and $b_{i}^{\rm left}$ (blue) $\ (i=1,2)$ are kept positive for almost all time. Only $b_{2}^{\rm left}$ violates the constraint at around $180$\mathrm{s}$$ due to the relaxation of the associated constraint (59c) and various uncertainties of the environment and the system model. Nevertheless, we see from the movie that the collisions with the walls actually do not happen throughout the experiment owing to the margin set in the definition of the environment. The movie also shows that the USVs successfully take avoidance actions when they are close to the wall, while giving up the path following. The USVs also tend not to go to the top-left and bottom-right areas, and such areas tend to have high importance indices. This is because the USVs have difficulty reaching these areas due to prioritizing $b_{i}^{\rm right}\geq 0$ over $b_{i}^{\rm left}\geq 0$ in (59). Actually, such phenomena cannot be observed in the simulation.
@@
V-E Comparative Evaluation with Baseline Algorithm
We finally conduct a comparative study between the present online coverage path generator and a lawnmower pattern algorithm. The main objective of the study is to reveal the benefits of the real-time feedback in the path generation. To this end, we take a control methodology that decouples the path planning and path following as a baseline algorithm.
Figure 19: Preset paths and the initial positions of the USV fleets.
Since existing path planning methods only generate finite-length paths, a modification is introduced for a fair comparison with the present method. To this end, the field is first divided into two areas as shown in Fig. 19. The path to be followed by each USV $i$ is generated to cover the assigned area using the Dubins Coverage with Area Clustering (DCAC) algorithm in [37]. The width of the adjacent stripes was set to $0.4$\mathrm{m}$$ based on the sensing radius $\sigma$ and the minimum turning radius $r_{\mathrm{min}}$. To move each USV persistently cover the assigned area, endpoints are connected forming the closed paths for persistent monitoring of Fig. 19.
These paths are discretized by $0.2$\mathrm{m}$$ and the waypoints $\varpi_{i}^{1},\varpi_{i}^{2},\cdots,\varpi_{i}^{\ell}$ are assigned at the endpoints of the path segments, with $\varpi_{i}^{\ell+1}=\varpi_{i}^{1}$. The closest segment is then selected as the initial target path segment for a path following controller based on the Line-of-Sight guidance algorithm [30]. Given the target path segment $\varpi_{i}^{k}\varpi_{i}^{k+1}$, the reference angle for USV $i$ is determined as follows:
$\theta_{i}^{\mathrm{ref}}=\vartheta_{\varpi,i}-\arctan\left(\frac{\mathbf{e}_{2}^{\top}R_{\vartheta_{\varpi,i}}p_{i}}{\Delta}\right),$ (60)
where $\vartheta_{\varpi,i}=\operatorname*{atan2}2(\mathbf{e}_{2}^{\top}(\varpi_{i}^{k+1}-\varpi_{i}^{k}),\mathbf{e}_{1}^{\top}(\varpi_{i}^{k+1}-\varpi_{i}^{k}))$ and $\Delta=0.5$. The control input is then determined by the PI controller
$u_{i}=-\left(1.3+\frac{0.14}{s}\right)(\theta_{i}^{\mathrm{ref}}-\theta_{i}).$ (61)
When the position of the USV gets closer than a threshold $D=0.3$\mathrm{m}$$ to the end of the target path segment $\varpi_{i}^{k+1}$, $\varpi_{i}^{k+1}\varpi_{i}^{k+2}$ is selected as a next target path segment.
The importance indices are updated by (2), based on the positions of the USVs driven by the above control strategy. The parameters of the update law (2) are selected as in Section V-D. However, notice that the baseline algorithm inherently cannot take into account the decay rate of the indices in designing the paths.
The video of the baseline algorithm experiment is available at https://youtu.be/ACtlfoO1MlU?si=uVJx-2Ql_PtJFcsM. The path is set up so that USVs do not collide with the pool wall. This is especially the case when the USV follows a path segment connecting the end and beginning points of the path generated by the DCAC algorithm. It takes about $55$\mathrm{s}$$ for a USV to complete one lap. When the USV follows the straight line part, the observation area tends to cover more points with higher importance indices. When it follows the curves, it tends to cover points that have been recently observed. The comparison of the sum of importance indices between the baseline algorithm and the proposed method is shown in Fig. 20. In both cases, the USVs pass through unvisited points at the beginning, so that there is no significant difference in the first $10$\mathrm{s}$$. The steep declines continue in the first $50$\mathrm{s}$$ in both methods but the coverage path generation achieves lower values than the lawnmower pattern. After the rapid decline, the values are stabilized, albeit with some oscillations. The results demonstrate that our proposed method achieves more efficient coverage by reducing the sum of importance indices more effectively than the baseline algorithm.
In addition, the present method is advantageous over the conventional path planning methods in the following aspects. First, the proposed path generator explicitly reflects the decay/increase rate of the importance indices which correspond to the required amount of the sensing data and the required timing of resampling. Second, the generator can compensate for the coverage holes caused e.g. by disturbances due to the real-time feedback of the current importance indices. Third, the generator has flexibility in adding or deleting USVs in the middle of the operation, since the area partitions are updated in real-time.
Figure 20: Time evolution of the sum of importance indices $\sum_{j=1}^{m}\phi_{j}$ of the proposed method in Section V-D (blue) and the baseline algorithm (red).
@@ VI Conclusion
Appropriately adjusting the vehicle speed together with the update rate of the path according to the dynamics of the environment is left as future work. Future work should also be directed towards an extension to a more generic class of paths and field experiments.
@@ Appendix A Quadratic Program for Circular Path Generation
Given a turning direction $\mathrm{X}_{i}\in\mathcal{X}$, we have
@@
$\mathcal{X}_{\mathrm{c},i}^{\epsilon}(r_{i})=\left\{\mathrm{X}_{i}\in\mathcal{X}\left|\ \left|\left(\max_{\mathrm{X}_{i}\in\mathcal{X}}I^{{\rm X}_{i}}_{\mathrm{c},i}(r_{i})\right)-I^{{\rm X}_{i}}_{\mathrm{c},i}(r_{i})\right|\leq\epsilon\right.\right\},$ (63)
@@ Appendix B Quadratic Program for Elliptic Path Generation
$\mathcal{X}_{\mathrm{e},i}^{\epsilon}(s_{i})$ =$=\left\{\mathrm{X}_{i}\in\mathcal{X}\left|\ \left|\left(\max_{\mathrm{X}_{i}\in\mathcal{X}}I^{{\rm X}_{i}}_{\mathrm{e},i}(s_{i})\right)-I^{{\rm X}_{i}}_{\mathrm{e},i}(s_{i})\right|\leq\epsilon\right.\right\},$